# Analysis and Simulation of a Fluid-Heat System in a Thin, Rough Layer in Contact With a Solid Bulk Domain

### v2

@@ 3.2 Analysis of the micro model
00 0 $=\lim_{k\to\infty}\left[(\mu(\psi_{k})\nabla u_{k},\nabla v)_{S\times\Omega^{f}_{\varepsilon}}-(p_{k},\operatorname{div}{(v)})_{S\times\Omega^{f}_{\varepsilon}}\right]=(\mu(\psi)\nabla u,\nabla v)_{S\times\Omega^{f}_{\varepsilon}}-(p,\operatorname{div}{(v)})_{S\times\Omega^{f}_{\varepsilon}},$
for all $v\in H^{1}_{0}(\Omega^{f}_{\varepsilon})^{d}$. By localization in time, we see that, for almost all $t\in S$, $(u,p)$ is the solution to the Stokes problem (3.2) in $V_{\varepsilon}\times Q_{\varepsilon}$ corresponding to $\psi$. Because the solution is unique, the whole sequence weakly converges to the same limit. ∎
@@
Given the assumptions, the sequences $\{\theta_{k}\}_{k\in\mathbb{N}}:=\{\mathbb{T}_{\varepsilon}(u_{k})\}_{k\in\mathbb{N}}\subset L^{2}(S;\Theta_{\varepsilon})$ and $\{\partial_{t}\theta_{k}\}_{k\in\mathbb{N}}\subset L^{2}(S;\Theta_{\varepsilon}^{\prime})$ are bounded uniformly in $k\in\mathbb{N}$ (see Lemma 4). Therefore, there is a subsequence, still indexed with $k$, and a weak limit $\theta\in L^{2}(S;\Theta_{\varepsilon})$ with $\partial_{t}\theta\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ such that $\theta_{k}\rightharpoonup\theta$ in $L^{2}(S;\Theta_{\varepsilon})$ and $\partial_{t}\theta_{k}\rightharpoonup\partial_{t}\theta$ in $L^{2}(S;\Theta_{\varepsilon}^{\prime})$.
@@
It remains to be shown that $\theta=\mathbb{T}_{\varepsilon}(u)$. Looking at the weak form Eq. 3.1, we can immediately pass to the limit in most terms with only the convection term as a potential issue. Now, since $\theta_{k}\to\theta$ strongly in $L^{2}(S\times\Omega_{\varepsilon}^{f})$ and $u_{k}\rightharpoonup u$ weakly in $L^{2}(S\times\Omega_{\varepsilon}^{f})^{d}$, we have $u_{k}\theta_{k}\rightharpoonup u\theta$ weakly in $L^{1}(S\times\Omega_{\varepsilon}^{f})^{d}$. Via the embedding $H^{1}(\Omega^{f}_{\varepsilon})\hookrightarrow L^{4}(\Omega^{f}_{\varepsilon})$ and $\sup_{k\in\mathbb{N}}\|u_{k}\|_{V_{\varepsilon}}<\infty$, we also have
@@
The operator $\mathbb{F}_{\varepsilon}$ defined in Eq. 3.10 is continuous from $L^{2}(S\times\Omega)$ to $L^{2}(S\times\Omega)$.
@@
Let Assumptions (A1)-(A5) be satisfied. Then the operator $\mathbb{F}_{\varepsilon}$ defined in Eq. 3.10 has a fixed point in $L^{2}(S;\Theta_{\varepsilon})$.
@@
We observe that $L^{2}(S;\Theta_{\varepsilon})\subset L^{2}(S\times\Omega)$ and $\mathbb{F}_{\varepsilon}[L^{2}(S\times\Omega)]\subset L^{2}(S;\Theta_{\varepsilon})$ by Lemma 4 and 7. Furthermore, $\partial_{t}\psi\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ for all $\psi\in\mathbb{F}_{\varepsilon}[L^{2}(S\times\Omega)]$ with uniform bounds from Lemma 4 and 6 which are independent of $\psi$. As in the proof of Lemma 9, we can use the Aubin-Lions lemma to conclude that $\mathbb{F}_{\varepsilon}[L^{2}(S\times\Omega)]$ is compact in $L^{2}(S\times\Omega)$. Finally, $\mathbb{F}_{\varepsilon}$ is $L^{2}(S\times\Omega)$-continuous (Corollary 1). Therefore, we apply Schauder’s fixed point theorem and obtain the existence of at least one fixed point of the operator $\mathbb{F}_{\varepsilon}$ in $L^{2}(S;\Theta)$. ∎
The previous theorem ensures that, under the given assumptions, there exists a weak solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})$ of the problem (2.1) for all $\varepsilon>0$. By construction of the operator $\mathbb{F}_{\varepsilon}$ the fixed point also fulfills the estimates of Lemma 4 and Lemma 7. To ensure uniqueness, further assumptions on the regularity of the solution are required, as denoted in the following theorem.
@@
Let Assumptions (A1)(A5) be satisfied. If there is a solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ of the coupled system Eq. 3.1 and (3.2), such that
  • if $d=2:$ $\nabla\theta_{\varepsilon}\in L^{\infty}(S;L^{1+\delta}(\Omega_{\varepsilon}^{f})^{2})$ and $\nabla u_{\varepsilon}\in L^{\infty}(S;L^{2+\delta}(\Omega_{\varepsilon}^{f})^{2\times 2})$,
  • if $d=3:$ $\nabla\theta_{\varepsilon}\in L^{\infty}(S;L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})^{3})$ and $\nabla u_{\varepsilon}\in L^{\infty}(S;L^{3+\delta}(\Omega_{\varepsilon}^{f})^{3\times 3})$,
@@
We provide the proof in the case $d=3$, the other case follows by the same steps but with a better estimate in the Gagliardo-Nierenberg inequality. Suppose there are two sets of solutions $(\theta_{\varepsilon}^{i},u_{\varepsilon}^{i},p_{\varepsilon}^{i})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ ($i=1,2$) with their differences denoted by $(\overline{\theta_{\varepsilon}},\overline{u_{\varepsilon}},\overline{p_{\varepsilon}})$. Furthermore, let $(\theta_{\varepsilon}^{1},u_{\varepsilon}^{1})$ satisfy the additional gradient regularity. Using the weak formulations, it holds
@@
$\frac{1}{2}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+c_{g}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\frac{1}{2\varepsilon}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\frac{c_{f}}{\varepsilon}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\alpha\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Gamma_{\varepsilon})}^{2}+\frac{1}{\varepsilon}(u_{\varepsilon}^{2}\cdot\nabla\overline{\theta_{\varepsilon}},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}\\ \leq-\frac{1}{\varepsilon}(\overline{u_{\varepsilon}}\cdot\nabla\theta_{\varepsilon}^{1},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}},$
$\begin{split}&\underline{\mu}\|\nabla\overline{u_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}\leq((\mu({\theta_{\varepsilon}}^{1})-\mu({\theta_{\varepsilon}}^{2}))\nabla u_{\varepsilon}^{1},\nabla\overline{u_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}.\end{split}$
Take $r=\nicefrac{{2(3+\delta)}}{{1+\delta}}$, then $1=\nicefrac{{1}}{{2}}+\nicefrac{{1}}{{r}}+\nicefrac{{1}}{{3+\delta}}$. Using the Lipschitz property of the viscosity and Hölder’s and the Poincaré inequality, we are led to
$\|\overline{u_{\varepsilon}}\|_{H^{1}(\Omega_{\varepsilon}^{f})}^{2}\leq C\|\overline{\theta_{\varepsilon}}\|_{L^{r}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}.$ (3.11)
@@
via Eq. 3.5 resulting in (for $q=\nicefrac{{18+12\delta}}{{3+10\delta}}$)
$\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\llbracket\overline{\theta_{\varepsilon}}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}^{2}\\ \leq C_{\varepsilon}\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{L^{q}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{{L^{\nicefrac{{3}}{{2}}+\delta}}(\Omega_{\varepsilon}^{f})}.$ (3.12)
Summing Eqs. 3.11 and 3.12 and integrating over time, we get
$\|\overline{\theta_{\varepsilon}}(t)\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\overline{\theta_{\varepsilon}}(t)\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\int_{0}^{t}\left(\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\llbracket\overline{\theta_{\varepsilon}}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}^{2}+\|\overline{u_{\varepsilon}}\|_{H^{1}(\Omega_{\varepsilon}^{f})}^{2}\right)\,\mathrm{d}\tau\\ \leq C_{\varepsilon}\int_{0}^{t}\left(\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{L^{q}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{{L^{\nicefrac{{3}}{{2}}+\delta}}(\Omega_{\varepsilon}^{f})}+\|\overline{\theta_{\varepsilon}}\|_{{L^{r}}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}\right)\,\mathrm{d}\tau.$
Focusing on the second term on the right-hand side, we have
@@
The Gagliardo-Nierenberg interpolation inequality with $\lambda=\nicefrac{{3}}{{3+\delta}}$, yields
@@
Using $\nabla u_{\varepsilon}^{1}\in L^{\infty}(S;L^{3}(\Omega_{\varepsilon}^{f}))^{d}$ and the scaled Young’s inequality leads to
@@
For the first term on the right-hand side, we use $\nabla\theta_{\varepsilon}^{1}\in L^{\infty}(S;{L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})})$ to estimate
@@
which (again using Young’s inequality and the Gagliardo-Nierenberg interpolation) gives

### v3

@@ 3.2 Well-posedness and a prior estimates
$0$ $=\lim_{k\to\infty}\left[(\mu(\psi_{k})\nabla u_{k},\nabla v)_{S\times\Omega^{f}_{\varepsilon}}-(p_{k},\operatorname{div}{(v)})_{S\times\Omega^{f}_{\varepsilon}}\right]=(\mu(\psi)\nabla u,\nabla v)_{S\times\Omega^{f}_{\varepsilon}}-(p,\operatorname{div}{(v)})_{S\times\Omega^{f}_{\varepsilon}},$
for all $v\in H^{1}_{0}(\Omega^{f}_{\varepsilon})^{d}$. By localization in time, we see that, for almost all $t\in S$, $(u,p)$ is the solution to the Stokes problem (8) in $V_{\varepsilon}\times Q_{\varepsilon}$ corresponding to $\psi$. Because the solution is unique, the whole sequence weakly converges to the same limit. ∎
@@
Given the assumptions, the sequences $\{\theta_{k}\}_{k\in\mathbb{N}}\coloneqq\{\mathbb{T}_{\varepsilon}(u_{k})\}_{k\in\mathbb{N}}\subset L^{2}(S;\Theta_{\varepsilon})$ and $\{\partial_{t}\theta_{k}\}_{k\in\mathbb{N}}\subset L^{2}(S;\Theta_{\varepsilon}^{\prime})$ are bounded uniformly in $k\in\mathbb{N}$ (see Lemma 5). Therefore, there is a subsequence, still indexed with $k$, and a weak limit $\theta\in L^{2}(S;\Theta_{\varepsilon})$ with $\partial_{t}\theta\in L^{2}(S;\Theta_{\varepsilon}^{\prime})$ such that $\theta_{k}\rightharpoonup\theta$ in $L^{2}(S;\Theta_{\varepsilon})$ and $\partial_{t}\theta_{k}\rightharpoonup\partial_{t}\theta$ in $L^{2}(S;\Theta_{\varepsilon}^{\prime})$.
@@
It remains to be shown that $\theta=\mathbb{T}_{\varepsilon}(u)$. Looking at the weak form Eq. 7, we can immediately pass to the limit in most terms with only the convection term as a potential issue. Now, since $\theta_{k}\to\theta$ strongly in $L^{2}(S\times\Omega_{\varepsilon}^{f})$ and $u_{k}\rightharpoonup u$ weakly in $L^{2}(S\times\Omega_{\varepsilon}^{f})^{d}$, we have $u_{k}\theta_{k}\rightharpoonup u\theta$ weakly in $L^{1}(S\times\Omega_{\varepsilon}^{f})^{d}$. Via the embedding $H^{1}(\Omega^{f}_{\varepsilon})\hookrightarrow L^{4}(\Omega^{f}_{\varepsilon})$ and $\sup_{k\in\mathbb{N}}\|u_{k}\|_{V_{\varepsilon}}<\infty$, we also have
@@
The operator $\mathbb{F}_{\varepsilon}$ defined in Eq. 16 is continuous from $L^{2}(S\times\Omega)$ to $L^{2}(S\times\Omega)$.
@@
Let Assumptions (A1)-(A5) be satisfied. Then, for all $\varepsilon>0$, the operator $\mathbb{F}_{\varepsilon}$ defined in Eq. 16 has a fixed point in $L^{2}(S;\Theta_{\varepsilon})$.
@@
By Lemma 5 and 7 there exists a $\varepsilon$-dependent $C_{\varepsilon}>0$ such that
$\|\mathbb{F}_{\varepsilon}(\psi)\|_{L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon}\cup\Omega^{f}_{\varepsilon}))}+\|\partial_{t}(\mathbb{F}_{\varepsilon}(\psi))\|_{L^{2}(S;\Theta_{\varepsilon}^{\prime})}\leq C_{\varepsilon},\quad\text{ for all }\psi\in L^{2}(S\times\Omega).$
Define the set
$M\coloneqq\left\{\psi\in L^{2}(S;\Theta_{\varepsilon}):\partial_{t}\psi\in L^{2}(S;\Theta_{\varepsilon}^{\prime}),\|\psi\|_{L^{2}(S;H^{1}(\Omega^{s}_{\varepsilon}\cup\Omega^{f}_{\varepsilon}))}+\|\partial_{t}\psi\|_{L^{2}(S;\Theta_{\varepsilon}^{\prime})}\leq C_{\varepsilon}\right\}.$
By construction, $M$ is closed, bounded, convex, and nonempty. Furthermore, by Lemma 5 and 8 it holds $\mathbb{F}_{\varepsilon}:M\to M$. We know that $\mathbb{F}_{\varepsilon}$ is $L^{2}(S\times\Omega)$-continuous (Corollary 1). As in the proof of Lemma 10, we can use the Aubin-Lions lemma to conclude that $M$, and also $\mathbb{F}_{\varepsilon}(M)$, are compact in $L^{2}(S\times\Omega)$. Therefore, we apply Schauder’s fixed point theorem [47, Theorem 1.C] and obtain the existence of at least one fixed point of the operator $\mathbb{F}_{\varepsilon}$ in $M$, and subsequently in $L^{2}(S;\Theta_{\varepsilon})$. ∎
The previous theorem ensures that, under the given assumptions, there exists a weak solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})$ of the problem (1) for all $\varepsilon>0$. By construction of the operator $\mathbb{F}_{\varepsilon}$ the fixed point also fulfills the estimates of Lemma 5 and Lemma 8. To ensure uniqueness, further assumptions on the regularity of the solution are required, as denoted in the following theorem.
@@
Let Assumptions (A1)(A5) be satisfied. If there is a solution $(\theta_{\varepsilon},u_{\varepsilon},p_{\varepsilon})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ of the coupled system Eq. 7 and (8), such that
  • if $d=2:$ $\nabla\theta_{\varepsilon}\in L^{\infty}(S;L^{1+\delta}(\Omega_{\varepsilon}^{f}))^{2}$ and $\nabla u_{\varepsilon}\in L^{\infty}(S;L^{2+\delta}(\Omega_{\varepsilon}^{f}))^{2\times 2}$,
  • if $d=3:$ $\nabla\theta_{\varepsilon}\in L^{\infty}(S;L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f}))^{3}$ and $\nabla u_{\varepsilon}\in L^{\infty}(S;L^{3+\delta}(\Omega_{\varepsilon}^{f}))^{3\times 3}$,
@@
We provide the proof in the case $d=3$, the other case follows by the same steps but with a better estimate in the Gagliardo-Nirenberg inequality. Suppose there are two sets of solutions $(\theta_{\varepsilon}^{i},u_{\varepsilon}^{i},p_{\varepsilon}^{i})\in L^{2}(S;\Theta_{\varepsilon})\times V_{\varepsilon}\times Q_{\varepsilon}$ ($i=1,2$) with their differences denoted by $(\overline{\theta_{\varepsilon}},\overline{u_{\varepsilon}},\overline{p_{\varepsilon}})$. Furthermore, let $(\theta_{\varepsilon}^{1},u_{\varepsilon}^{1})$ satisfy the additional gradient regularity. Using the weak formulations, it holds
@@
$\frac{1}{2}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\kappa_{g}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\frac{1}{2\varepsilon}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\frac{\kappa_{f}}{\varepsilon}\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\alpha\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Gamma_{\varepsilon})}^{2}+\frac{1}{\varepsilon}(u_{\varepsilon}^{2}\cdot\nabla\overline{\theta_{\varepsilon}},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}\\ =-\frac{1}{\varepsilon}(\overline{u_{\varepsilon}}\cdot\nabla\theta_{\varepsilon}^{1},\overline{\theta_{\varepsilon}})_{\Omega^{f}_{\varepsilon}},$
$\begin{split}&\underline{\mu}\|\nabla\overline{u_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}=((\mu({\theta_{\varepsilon}}^{1})-\mu({\theta_{\varepsilon}}^{2}))\nabla u_{\varepsilon}^{1},\nabla\overline{u_{\varepsilon}})_{\Omega^{f}_{\varepsilon}}.\end{split}$
Take $r=\nicefrac{{2(3+\delta)}}{{1+\delta}}$, then $1=\nicefrac{{1}}{{2}}+\nicefrac{{1}}{{r}}+\nicefrac{{1}}{{3+\delta}}$. Using the Lipschitz property of the viscosity and the Hölder and the Poincaré inequalities, we are led to
$\|\overline{u_{\varepsilon}}\|_{H^{1}(\Omega_{\varepsilon}^{f})}^{2}\leq C\|\overline{\theta_{\varepsilon}}\|_{L^{r}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}.$ (17)
@@
via Eq. 11, resulting in (for $q=\nicefrac{{18+12\delta}}{{3+10\delta}}$)
$\begin{split}\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\frac{\operatorname{d}}{\operatorname{d}t}\|\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\llbracket\overline{\theta_{\varepsilon}}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}^{2}\\ \leq C_{\varepsilon}\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{L^{q}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{{L^{\nicefrac{{3}}{{2}}+\delta}}(\Omega_{\varepsilon}^{f})}.\end{split}$ (18)
Summing Eqs. 17 and 18 and integrating over time, we get
$\begin{split}\|\overline{\theta_{\varepsilon}}(t)\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\overline{\theta_{\varepsilon}}(t)\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\int_{0}^{t}\left(\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{s})}^{2}+\|\nabla\overline{\theta_{\varepsilon}}\|_{L^{2}(\Omega_{\varepsilon}^{f})}^{2}+\|\llbracket\overline{\theta_{\varepsilon}}\rrbracket\|_{L^{2}(\Gamma_{\varepsilon})}^{2}+\|\overline{u_{\varepsilon}}\|_{H^{1}(\Omega_{\varepsilon}^{f})}^{2}\right)\,\mathrm{d}\tau\\ \leq C_{\varepsilon}\int_{0}^{t}\left(\|\overline{u_{\varepsilon}}\|_{L^{6}(\Omega_{\varepsilon}^{f})}\|\overline{\theta_{\varepsilon}}\|_{L^{q}(\Omega_{\varepsilon}^{f})}\|\nabla{\theta_{\varepsilon}}^{1}\|_{{L^{\nicefrac{{3}}{{2}}+\delta}}(\Omega_{\varepsilon}^{f})}+\|\overline{\theta_{\varepsilon}}\|_{{L^{r}}(\Omega_{\varepsilon}^{f})}^{2}\|\nabla u_{\varepsilon}^{1}\|_{{L^{3+\delta}}(\Omega_{\varepsilon}^{f})}^{2}\right)\,\mathrm{d}\tau.\end{split}$ (19)
Focusing on the second term on the right-hand side of Equation 19, we have
@@
The Gagliardo-Nirenberg interpolation inequality with $\lambda=\nicefrac{{3}}{{3+\delta}}$, yields
@@
Using $\nabla u_{\varepsilon}^{1}\in L^{\infty}(S;L^{3+\delta}(\Omega_{\varepsilon}^{f}))^{d}$ and the scaled Young’s inequality leads to
@@
For the first term on the right-hand side of Equation 19, we use $\nabla\theta_{\varepsilon}^{1}\in L^{\infty}(S;{L^{\nicefrac{{3}}{{2}}+\delta}(\Omega_{\varepsilon}^{f})})$ to estimate
@@
which (again using Young’s inequality and the Gagliardo-Nirenberg interpolation) gives